OpenStudy (anonymous):

solve x+y+z=8 x-y-z=4 -x+y-z=8

OpenStudy (anonymous):

x+y+z=8 x-y-z=4 -x+y-z=8 Process of elimination. x+y+z=8 x-y-z=4 Add 2x=12 Divide x=6 ~~~~~~~~~~~~~~~ (x-y-z=4) times -1 -x+y+z=-4 -x+y-z=8 Add -2x+2y=4 ~~~~~~~~~~~~~~~~ (x=6) times 2 2x=12 -2x+2y=4 Add 2y=16 Divide y=8 ~~~~~~~~~~~~~~ Substitute answers into original equation. 6+8+z=8 14+z=8 Subtract z=-6 ~~~~~~~~~~~~~~~~~~~ Substitute all variable answers in to all equations to see if correct. 6+8+(-6)=8 14-6=8 8=8 Correct! 6-8-(-6)=4 -2-(-6)=4 -2+6=4 4=4 Correct! -6+8-(-6)=8 2-(-6)=8 2+6=8 8=8 Correct! ~~~~~~~~~~~~~~~~~~ So x=6 y=8 z=-6

OpenStudy (anonymous):

Get it?